# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

### v1

@@ Appendix A Related work
We follow standard expositions in Baader-Nipkow and Terese [BaaderNipkow98, Terese03]. Termination under duplication relies on DM multiset orders [DershowitzManna79] and (recursive) path orders [Dershowitz87]. Newman’s Lemma is used in its classic ARS form [Newman42]. For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [Goodstein44, KirbyParis82, Buchholz86]. For fixed-target reachability complexity and modularity results, see [BaaderGiesl24, CaronCoquide94]; for undecidability at shallow depth in non-linear systems, see [Mitsuhashi06].
@@ Appendix B Conclusion
We gave an operator-only kernel and proved strong normalization and confluence for the guarded safe subrelation, yielding a certified normalizer and unique normal forms via Newman’s Lemma. We also established an explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$), blocking confluence for the full relation; this motivates the designed SafeStep fragment. A decidability result explains why single-level Gödel encodings clash with global termination. We proved a simple impossibility for additive bumps under duplication and certified DM/MPO orientations. Future work: complete a stratified meta-layer; develop arithmetic encodings within KO7; and map the exact boundary between internally definable measures and external ordinal strength.

### v2

@@ Appendix A Related work
This work follows standard expositions in Baader-Nipkow and Terese [2, 11]. Termination under duplication relies on DM multiset orders [5] and (recursive) path orders [6]. Newman’s Lemma is used in its classic ARS form [10]. For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [7, 8, 3]. For fixed-target reachability complexity and modularity results, see [1, 4]; for undecidability at shallow depth in non-linear systems, see [9].
@@ Appendix B Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation. These results yield a certified normalizer and unique normal forms via Newman’s Lemma. An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment. The decidability result explains the clash between single-level Gödel encodings and global termination. The impossibility of additive bumps under duplication is formally verified. Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.